[ two buses leave Orlando at the same time and head in opposite directions. the northbound bus averaging 55 mph while the southbound bus is averaging 45 mph. after how many hours will the buses [be] 450 miles apart? ]
Answer by TimothyLamb(4379)

since we're only concerned with the relative distance between the two buses, we can simplify the problem to a single bus moving at the average relative speed of the two buses.
---
since the buses are traveling in exactly opposite directions the vector math for speed is very simple:
s = 55 + 45 = 100
---
speed of the two buses relative to each other: 100 mph
---
s = d / t
t = d / s
---
time:
t = d / s
t = 450 / 100
t = 4.5
---
Answer:
after 4.5 hours the two buses will be 450 miles apart
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php